• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 15:51:56 on January28,2002

    Entry number 81104

    keyword=Day shift summary

    08:00 In Controlled access to fix spectrometer motion. -->M. Stevens
    A. Glamazdin inserted post radiator. Last run without post radiator is 1801.
    10:30 Spectrometer work completed. Beam permit.
    10:45 Beam on carbon target
    11:20 Move target to H2. two reboots. Christopher Keith runs some diagnostics.
    11:50 Target in positon. Beam to hall
    12:30 MCC fixing minor problems. Beam down
    01:07 Beam back to hall
    01:55 gas alarm. Also scintillator high voltages out of range during run 1806
    02:15 Accessed gas shed to override low pressure alarm.
    03:00 ROC4 not responding. 100% deadtime. Reboot.
    03:12 resume data taking.
    03:42 completed kinematic 4E. Call for access.